### v1

# Global convergence of a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems

Abstract: We propose a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems (MFQNMO). The method employs a common BFGS matrix to approximate the Hessian matrix of all objective functions in each iteration. This matrix is updated using function and gradient information from the previous step. Our analysis confirms the convergence of the method without relying on convexity assumptions. We also establish a local superlinear convergence rate for MFQNMO under mild conditions and validate its effectiveness through experiments on both nonconvex and convex test problems.
Keywords: Multiobjective optimization, quasi-Newton methods, nonconvex, Wolfe line search, global convergence.
AMS subject classifications:
1 Introduction
In this paper, we consider the following unconstrained multiobjective optimization problems (MOP):
$\min_{x\in R^{n}}f(x)=\left(f_{1}\left(x\right),f_{2}\left(x\right),\ldots,f_{m}\left(x\right)\right)^{T},$ (MOP)
where $f_{i}\left(x\right){:}R^{n}\to R$ is continuously differentiable function, for any $i\in[m]$. Multiobjective optimization is widely used in many research areas, which can be found in engineering [1, 2], environmental analysis [3, 4], machine learning [5, 6], etc.
In 2000, Fliege and Svaiter proposed the steepest descent methods for multiobjective optimization [7]. Building on their pioneering work, researchers have conducted extensive studies on the first-order descent algorithm for solving (MOP), (see e.g. [11, 12, 13, 14]). The first-order algorithm is often favored for its lower cost compared to the second-order method, as it avoids the computational burden of determining the second-order Hessian matrix. However, the trade-off is a typically slower convergence rate, particularly evident in scenarios involving ill-conditioned problems.
The ill-conditioned problem may exhibit large curvature variations at certain points, leading to highly unstable curvature values. The Newton method [8] is effective in such cases as it captures the local curvature of functions and adjusts the direction more efficiently, leading to faster rate compared to first-order algorithms. Fliege et al. [8] proved the superlinear convergence rate of the Newton method. However, despite its efficiency, the computational cost per iteration is notably high, particularly for large-scale problems.
To balance the computational cost and the effectiveness, the quasi-Newton method is considered. Different from single objective optimization, the quasi-Newton method [9] uses different matrices to approximate the Hessian matrix of individual objective functions in each iteration to capture the curvature information. Numerous variants of the quasi-Newton method have been proposed. Lai et al. [16] proposed a quasi-Newton method for constructing approximate matrices using q-derivate. In [17], the global convergence of the quasi-Newton method is established by the average nonmonotone line search. Morovati et al. [18] extended both the self-scaling BFGS (SS-BFGS) method and the Huang BFGS (H-BGFS) method for multiobjective optimization problems. Prudente and Souza [19], operating under the assumption of no convexity, proposed a modified BFGS formula. They demonstrated that their algorithm is well-defined under the Wolf line search without convexity assumption. Different from the case of single-objective, the computational cost per iteration still be high when the number of objective functions increases for multiobjective optimization problems. To address this, Ansary and Panda [15] utilized a common matrix to approximate all Hessian matrices in each iteration. Lapucci and Mansueto [21] extended the limited memory Quasi-Newton approach for MOPs (L-BFGS). However, all the improved algorithms mentioned above require convexity conditions to ensure global convergence. In [10, 20], the global convergence can be guaranteed with Wolf line search in the nonconvex case. The approximate Secant equation in reference [20] depends on some parameters, and updating the BFGS formula under strong convexity conditions does not always ensure a satisfactory approximation of the Secant equation.
For MOPs, there are relatively few studies on quasi-Newton methods in nonconvex cases. This raises the question of whether alternative methods exist that can balance between algorithmic convergence and per-iteration cost for nonconvex MOPs. In this paper, inspired by [20, 15], we proposed a modified BFGS-type method based on function information for nonconvex MOPs (MFQNMO). By utilizing both function and gradient information from the previous iterations, we construct a common modified BFGS-type formula to approximate the Hessian matrices of all objective functions in each iteration. We verify that the modified BGFS-type matrix is positive definite for each iteration and establish a global convergence for MFQNMO without convexity assumption.
The structure of this paper is as follows: In sect. 2, we give some notations and definitions for our later use. Sect. 3 recalls several multiobjective gradient descent methods. In sect. 4, we introduce the proposed MFQNMO method and establish its convergence analysis. Numerical results are presented in sect. 5. At the end of the paper, we draw some conclusions.
2 Preliminaries
We give some notations used in this paper.
$\bullet$ $\left[m\right]=\left\{1,2,\cdots,m\right\}.$
$\bullet$ ${{\Delta}_{m}}=\left\{\lambda:\sum\limits_{i\in\left[m\right]}{{{\lambda}_{i}}=1},{{\lambda}_{i}}\geq 0,\ i\in\left[m\right]\right\}.$
[…]
$\bullet$ ${}Jf\left(x\right)\in{{R}^{m\times n}}$, $\nabla{{f}_{i}}\left(x\right)\in{{R}^{n}}$ ,$\nabla^{2}f_{i}(x)$ the Jacobian matrix, the gradient of ${{f}_{i}}$ and the Hessian matrix at $x$, respectively.
$\bullet$ $\mathcal{D}(x,d)\coloneqq\max\limits_{i\in[m]}\nabla f_{i}(x)^{\top}d.$
$\bullet$ For all $B\in R^{n\times n}$, $B\succ 0$ is equivalent to $B$ is positive definite.
$\bullet$ For $x,y\in R^{n}$, $\langle x,y\rangle_{B}:=x^{\top}By$ and $\|x\|_{B}:=\sqrt{\langle x,x\rangle_{B}}$.
We define partial order for our later use:
$A\preceq B\Leftrightarrow B-A\succeq 0,\quad A,B\in R^{n\times n}$
$f\left(y\right)\leq f\left(x\right)\Leftrightarrow{{f}_{i}}\left(x\right)-{{f}_{i}}\left(y\right)\geq 0,\ i\in\left[m\right].$
Definition 2.1.
[7] A point ${{x}^{*}}\in{{R}^{n}}$ is called the Pareto (weakly Pareto ) solution for MOP, if there exists no point $x\in{{R}^{n}}$ such that $f\left(x\right)\leq f\left({{x}^{*}}\right)$ and $f\left(x\right)\neq f\left({{x}^{*}}\right)~{}\left(f\left(x\right)<f\left({{x}^{*}}\right)\right)$.
Definition 2.2.
[7] A point ${{x}^{*}}\in{{R}^{n}}$ is called a Pareto critical point for MOP, if exists $i\in\left[m\right]$ such that $\left\langle\nabla{{f}_{i}}\left({{x}^{*}}\right),d\right\rangle\geq 0,$ for all $d\in{{R}^{n}}.$
Definition 2.3.
[7] A vector $d\in{{R}^{n}}$ is called descent direction for $f$ at ${x}^{*}$, if $\left\langle\nabla{{f}_{i}}\left({{x}^{*}}\right),d\right\rangle<0,$ for all $i\in\left[m\right]$.
Lemma 2.1.
[8, Theorem 3.1] The following statements hold:
(i) if $x^{\ast}$ is local weak Pareto optimal, then $x^{\ast}$ is a Pareto critical point for $f$;
(ii) if $f$ is convex and $x^{\ast}$ is Pareto critical for $f$, then $x^{\ast}$ is weak Pareto optimal;
(iii) if $f$ is strictly convex and $x^{\ast}$ is Pareto critical for $f$, then $x^{\ast}$ is Pareto optimal.
3 Gradient descent methods for MOPs.
In this section, we revisit the steepest descent method [7] and the quasi-Newton method [9] for MOPs.
3.1 The steepest descent method for MOPs.
For a given $x\in{{R}^{n}}$, consider the following subproblem (3.1):
$\underset{d\in{{R}^{n}}}{\mathop{\min}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}.$ (3.1)
Since $\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}$ is strongly convex for $i\in\left[m\right]$, then subproblem (3.1) has a unique minimizer. We denote that $\theta\left(x\right)$ and $d\left(x\right)$ are the optimal value and the optimal solution in problem (3.1), respectively. Therefore,
$d\left(x\right)=\underset{d\in{{R}^{n}}}{\mathop{\text{argmin}}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}},$
and
$\theta\left(x\right)=\underset{d\in{{R}^{n}}}{\mathop{\min}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}.$
The problem can be transformed into the following equivalent optimization problem (3.2), which is easier to solve.
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t+\frac{1}{2}\|d\|^{2}\\ \text{s.t.}&\langle\nabla f_{i}(x),d\rangle\leq t,\quad i\in[m].\end{array}$ (3.2)
According to the Slater condition, strong duality is established. The dual problem has the same minimum value as (3.2). We consider that the dual problem of problem (3.2):
$\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\max}}\,\underset{\left(t,d\right)\in R\times{{R}^{n}}}{\mathop{\min}}\,L\left(\lambda,\left(t,d\right)\right),$
where $L\left(\lambda,\left(t,d\right)\right)$ is a Lagrange function,
$L\left(\lambda,\left(t,d\right)\right)=t+\frac{1}{2}{{\left\|d\right\|}^{2}}+\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}}\left(\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle-t\right).$
According to the KKT condition, we can obtain the steepest descent direction:
$d_{SD}\left(x\right)=-\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}^{SD}}\left(x\right)\nabla{{f}_{i}}\left(x\right),$
where $\lambda_{SD}\left(x\right)={{\left({{\lambda}_{1}^{SD}}\left(x\right),{{\lambda}_{2}^{SD}}\left(x\right),\ldots,{{\lambda}_{m}^{SD}}\left(x\right)\right)}^{T}}$is the solution of the following dual problem
$-\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\min}}\,\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}}\nabla{{f}_{i}}\left(x\right)\right\|}^{2}},$ (3.3)
owing to strong duality, we have:
$\theta_{SD}\left(x\right)=-\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}^{SD}}\left(x\right)\nabla{{f}_{i}}\left(x\right)\right\|}^{2}}=-\frac{1}{2}{{\left\|d_{SD}\left(x\right)\right\|}^{2}}\leq 0,$
$\left\langle\nabla{{f}_{i}}\left(x\right),d_{SD}\left(x\right)\right\rangle\leq-{{\left\|d_{SD}\left(x\right)\right\|}^{2}}\leq 0.$
The following lemma describes the proposition of $\theta_{SD}\left(x\right)$ and $d_{SD}\left(x\right).$
Lemma 3.1.
[7, Proposition 3.1] For problem (3.1), $\theta_{SD}\left(x\right)$ and $d_{SD}\left(x\right)$ are the optimal value and solution, respectively. Then,
$(a)$ It can be conclude that the equivalent relation is as follows.
$(i)$ The point $x\in{{R}^{n}}$ is not Pareto critical;
$(ii)$ $d_{SD}\left(x\right)\neq 0$;
$(iii)$ $\theta_{SD}\left(x\right)<0$.
$(b)$ The function $d_{SD}\left(x\right)$ and $\theta_{SD}\left(x\right)$ are continuous, for all $x\in{{R}^{n}}.$
3.2 The quasi-Newton methods for MOPs (QNMO).
Quasi-Newton direction $d_{q}(x)$ in [9] is the optimal solution to the following subproblem:
$\min_{d\in R^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d,$ (3.4)
where $B_{i}(x)\in R^{n\times n}$, approximates the Hessian matrix $\nabla^{2}f_{i}(x)$ and $B_{i}(x)\succ 0$, for $i\in[m]$. $\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d$ is strongly convex, for $i\in[m]$. The subproblem 3.4 has a unique minimizer. We denote that $\theta_{q}\left(x\right)$ and $d_{q}\left(x\right)$ be the optimal value and solution in problem (3.4), respectively,
$d_{q}(x)=\mathop{\text{argmin}}_{d\in R^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d,$
and
$\theta_{q}(x)=\max_{i\in[m]}\nabla f_{i}(x)^{\top}d(x)+\frac{1}{2}d(x)^{\top}B_{i}(x)d(x).$
(3.4) is equivalent to the convex quadratic optimization problem:
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t\\ \text{s.t.}&\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d\leq t,\quad\forall i\in[m].\end{array}$
According to the KKT condition, we can obtain the quasi-Newton direction:
$d_{q}(x)=-\left[\sum_{i=1}^{m}\lambda_{i}^{q}(x)B_{i}(x)\right]^{-1}(\sum_{i=1}^{m}\lambda_{i}^{q}(x)\nabla f_{i}(x)),$ (3.5)
where $\lambda_{q}\left(x\right)={{\left({{\lambda}_{1}^{q}}\left(x\right),{{\lambda}_{2}^{q}}\left(x\right),\ldots,{{\lambda}_{m}^{q}}\left(x\right)\right)}^{T}}$is the solution of the following dual problem
$-\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\min}}\,\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}}\nabla{{f}_{i}}\left(x\right)\right\|^{2}_{\left[\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}{B_{i}(x)}}\right]^{-1}}}}.$
Owing to strong duality, we have
$\theta_{q}(x)=-\frac{1}{2}d_{q}(x)^{\top}\left[\sum_{i=1}^{m}\lambda_{i}^{q}(x)B_{i}(x)\right]d_{q}(x).$
3.3 The modified quasi-Newton methods for MOPs (MQNMO).
Even if quasi-Newton avoids calculating the second-order Hessian matrix, the per-iteration cost is still expensive. Recently, Ansary and Panda [15] proposed a modified quasi-Newton method (MQNMO) that utilized a single positive matrix to approximate all the Hessian matrices. The Modified quasi-Newton direction $d(x)$ is the optimal solution to the following subproblem 3.6:
$\min_{d\in R^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B(x)d,$ (3.6)
where $B(x)\in R^{n\times n}$ and $B(x)\succ 0$. Problem (3.6) is equivalent to the convex quadratic optimization problem as follows
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t\\ \text{s.t.}&\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B(x)d\leq t,\quad\forall i\in[m].\end{array}$
According to the KKT condition, we can obtain the modified quasi-Newton direction
$d(x)=-B(x)^{-1}\left(\sum_{i=1}^{m}\lambda_{i}(x)\nabla f_{i}(x)\right),$ (3.7)
and
$\theta(x)=-\frac{1}{2}d(x)^{\top}B(x)d(x),$ (3.8)
where $\lambda\left(x\right)={{\left({{\lambda}_{1}}\left(x\right),{{\lambda}_{2}}\left(x\right),\ldots,{{\lambda}_{m}}\left(x\right)\right)}^{T}}$is the solution of the following dual problem (3.9)
$-\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\min}}\,\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}}\nabla{{f}_{i}}\left(x\right)\right\|}_{B(x)^{-1}}^{2}}.$ (3.9)
According to the above analysis, we can get
$\mathcal{D}(x_{k},d_{k})\leq-\|d_{k}\|_{B_{k}}^{2}$ (3.10)
The following rank one DFP formula of generating $B_{k+1}$ is defined in [15] by
$B_{k+1}=(I-\frac{y_{k}s_{k}^{\top}}{s_{k}^{\top}y_{k}})B_{k}(I-\frac{s_{k}y_{k}^{\top}}{s_{k}^{\top}y_{k}})+\frac{y_{k}y_{k}^{\top}}{s_{k}^{\top}y_{k}},$ (3.11)
where $s_{k}=x_{k+1}-x_{k}$ and $y_{k}=\sum_{i=1}^{m}\lambda_{i}^{k}(\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})).$
4 MFQNMO:Modified BFGS method based on function information for MOPs.
In this section, we propose a modified BFGS method based on function information for MOPs and analyze its properties. The formal description of the algorithm is as follows.
Algorithm 1 MFQNMO
1:Choose starting point $x^{0}\in\mathbb{R}^{n}$ and parameters $\sigma_{1},\sigma_{2}\in(0,1).$ Let $k=0$ and $B_{0}\succ 0.$
2:Compute the solution ${{\lambda}_{k}}={{\left({{\lambda}_{1}}\left({{x}_{k}}\right),{{\lambda}_{2}}\left({{x}_{k}}\right),\ldots,{{\lambda}_{m}}\left({{x}_{k}}\right)\right)}^{T}}$ of problem (3.9) and ${{d}_{k}}$, $\theta_{k}$ as (3.7), (3.8).
3:If $\left|{{\theta}_{k}}\right|<\varepsilon$, return $x_{k}.$
4:Compute step size $\alpha_{k}>0$ satisfies
$F_{i}(x_{k}+\alpha_{k}d_{k})$ $\leq F_{i}(x_{k})+\rho\alpha_{k}\mathcal{D}(x_{k},d_{k}),\quad\forall i\in[m],$ (4.1)
$\mathcal{D}(x^{k}+\alpha_{k}d_{k},d_{k})$ $\geq\sigma\mathcal{D}(x_{k},d_{k}),$ (4.2)
5:Set $x_{k+1}=x_{k}+\alpha_{k}d_{k}$.
6:Compute
$\eta_{k}=\dfrac{y_{k}^{\top}s_{k}}{\|s_{k}\|^{2}},$ (4.3)
$m_{k}=\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1})).$ (4.4)
$\gamma_{k}=y_{k}+m_{k}s_{k}.$ (4.5)
where $s_{k}=x_{k+1}-x_{k}$ , $\mu_{i}^{k}=\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})$ and $y_{k}=\sum_{i=1}^{m}\lambda_{i}^{k}\mu_{i}^{k}$.
7:Update the BFGS-type matrices,
$B_{k+1}=B_{k}-\dfrac{B_{k}s_{k}s_{k}^{\top}B_{k}}{s_{k}^{\top}B_{k}s_{k}}+\dfrac{\gamma_{k}\gamma_{k}^{\top}}{\gamma_{k}^{\top}s_{k}},$ (4.6)
$B_{k+1}^{-1}=(I-\frac{s_{k}\gamma_{k}^{\top}}{s_{k}^{\top}\gamma_{k}})B_{k}^{-1}(I-\frac{\gamma_{k}s_{k}^{\top}}{s_{k}^{\top}\gamma_{k}})+\frac{s_{k}s_{k}^{\top}}{s_{k}^{\top}\gamma_{k}}.$ (4.7)
8:Set $k\leftarrow k+1$ and go to Step 1.
Remark 4.1.
In step 1, solving the descent direction through the dual problem of subproblem 3.9, it is noted that the initial BFGS matrix is selected as a positive definite matrix, usually chosen as the unit matrix. Step 2 is the termination criterion of the algorithm. In step 3, we search the step size through the Wolfe line search. In steps 5-6, we notice that the BFGS-type update formula (4.6) and (4.7) is different from [20]. We use the previous iterative function information to update the $m_{k}$ adaptively. It can note that
$B_{k+1}s_{k}=y_{k}+m_{k}s_{k}.$ (4.8)
The following result shows that Algorithm 1 is well-defined without convexity assumption.
Theorem 4.1.
The BFGS matrix (4.6) updated by Algorithm 1 is positive definite.
Proof.
At the k-th iteration, we only need to consider the case when $x_{k}$ is not the Pareto critical point. By the definition of $\gamma_{k}$, $\eta_{k}$ and $m_{k}$, we have
$\gamma_{k}^{\top}s_{k}$ $=y_{k}^{\top}s_{k}+m_{k}\|s_{k}\|^{2}$
$=\left(\frac{y_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}+m_{k}\right)\|s_{k}\|^{2}$
$=(\eta_{k}+m_{k})\|s_{k}\|^{2}$
$=\left[\eta_{k}+\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x_{k})-f_{i}(x_{k+1})\right)\right]\|s_{k}\|^{2}$
$\geq\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x_{k})-f_{i}(x_{k+1})\right)\|s_{k}\|^{2}$ (4.9)
$>0.$ (4.10)
Then for all nonzero vector $x\in R^{n}$,
$x^{\top}B_{k+1}x=\|x\|_{B_{k}}^{2}-\dfrac{\langle x,s_{k}\rangle_{B_{k}}^{2}}{\|s_{k}\|_{B_{k}}^{2}}+\dfrac{(x^{\top}\gamma_{k})^{2}}{\gamma_{k}^{\top}s_{k}},$ (4.11)
it follows by the Cauchy-Schwarz inequality and (4.10) that
$x^{\top}B_{k+1}x\geq\dfrac{(x^{\top}\gamma_{k})^{2}}{\gamma_{k}^{\top}s_{k}}\geq 0.$ (4.12)
Inequality 4.12 holds as an equality when $x^{\top}\gamma_{k}=0$ and $x=cs^{k}$, where $c$ is a constant. It follows that $cs_{k}^{\top}\gamma_{k}=0$. By (4.10), we can obtain that $c=0$, which contradicts the fact that $x$ is a nonzero vector.
Assumption 4.1.
For any $z\in R^{n}$, the level set $\left\{x:f(x)\leq f(z)\right\}$ is a bounded set.
Next, we show that the Wolf line search in Algorithm 1 can terminate at a finite step.
Proposition 4.1.
Suppose Assumption 4.1 holds and $d_{k}$ is the descent direction. Then there exists an interval of values $[\alpha_{1},\alpha_{2}]$, with $0<\alpha_{1}<\alpha_{2}$, such that for all $\alpha\in[\alpha_{1},\alpha_{2}]$, inequalities (4.1) and (4.2) hold.
Proof.
The proof is similar to that in [21, Proposition 2], we omit it here. ∎
Lemma 4.1.
Let $\theta(x)$ and $d(x)$ be the optimal value and solution in problem (3.6). Then we have:
$(a)$ Vector $d(x)\neq 0$ is a descent direction.
$(b)$ The following conditions are equivalent:
$(i)$ The point $x\in{{R}^{n}}$ is not Pareto critical;
$(ii)$ $d(x)\neq 0$;
$(iii)$ $\theta(x)<0$.
$(c)$ If there exists $K\subset N$ such that $x_{k}\stackrel{{\scriptstyle K}}{{\longrightarrow}}x^{*}$, $d_{k}\stackrel{{\scriptstyle K}}{{\longrightarrow}}0$, then $x^{*}$ is a Pareto critical point.
Proof.
These assertions can be derived by using the same proof demonstrated in the [32, Proposition 4.1]. ∎
4.1 Global convergence
In this section, we present the global convergence analysis of Algorithm 1 without assuming convexity. First, we require the following assumptions.
Assumption 4.2.
(i) For $i\in[m]$, $\nabla f_{i}(x)$ is Lipschitz continuous with constant $L_{i}.$
(ii)For any $x\in R^{n}$, there exists constant $a$ and $b$ such that
$aI_{n}\preceq B(x)\preceq bI_{n}.$
In the subsection, the following properties imply that under the Wolfe line search, the objective function will be significantly reduced along the direction generated by Algorithm 1.
Lemma 4.2.
Let $B_{k}$ be generated by the BFGS update formula (4.6) in Algorithm 1. If there exists constants $c_{1}$ and $c_{2}$ such that $\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq c_{1}>0$, $\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq c_{2}$. Then, for any $p\in(0,1)$, there exists a constant $\delta>0$ such that, for any $k\in N$,
$\cos\beta_{j}=\frac{s_{j}^{\top}B_{j}s_{j}}{\|s_{j}\|\|B_{j}s_{j}\|}\geq\delta,$ (4.13)
holds for at least $\lceil pk\rceil$ values of $j\in[1,k].$
Proof.
The proof is similar to that in [22, Theorem 2.1], we omit it here. ∎
Proposition 4.2.
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Suppose that Assumption 4.2 holds. Then, there exists constants $c_{1}$ and $c_{2}$ such that
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq c_{1}\|d_{k}\|_{B_{k}}^{2},$ (4.14)
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq\frac{c_{2}}{\|d_{k}\|_{B_{k}}^{2}}.$ (4.15)
Proof.
First, inequality (4.14) can be deduced from Wolfe line search, (3.10) and (4.9), i.e.,
$f_{i}(x_{k})-f_{i}(x_{k+1})\geq-\sigma_{1}\alpha_{k}\mathcal{D}(x_{k},d_{k})\geq\sigma_{1}\alpha_{k}\|d_{k}\|_{B_{k}}^{2},$
$\gamma_{k}^{\top}s_{k}\geq\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x^{k})-f_{i}(x^{k+1})\right)\|s_{k}\|^{2}.$
We denote $c_{1}=\sigma_{1}\alpha_{1}>0$, it follows by Proposition 4.1 that (4.14) holds. Next, we consider that inequality (4.15). By the definition of the 4.5,
$\|\gamma_{k}\|$ $\leq\|y_{k}\|+m_{k}\|s_{k}\|$
$\leq\left(\frac{\|y_{k}\|}{\|s_{k}\|}+m_{k}\right)\|s_{k}\|$
$\leq\left[\frac{\|y_{k}\|}{\|s_{k}\|}+\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x^{k})-f_{i}(x^{k+1})\right)\right]\|s_{k}\|.$ (4.16)
Since $\nabla f_{i}(x)$ is Lipschitz continuous with constant $L_{i}$, for $i\in[m]$, we can get
$\frac{\|y_{k}\|}{\|s_{k}\|}\leq\sum_{i\in[m]}\lambda_{i}^{k}\frac{\|\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})\|}{\|x_{k+1}-x_{k}\|}\leq\max_{i\in[m]}L_{i}.$
By the Cauchy-Schwarz inequality,
$\frac{|\langle y_{k},s_{k}\rangle|}{\|s_{k}\|^{2}}\leq\frac{\|y_{k}\|}{\|s_{k}\|}.$
Therefore,
$\frac{\|y_{k}\|}{\|s_{k}\|}+\max\{-\eta_{k},0\}\leq\frac{2\|y_{k}\|}{\|s_{k}\|}\leq 2\max_{i\in[m]}L_{i}.$ (4.17)
Assumption 4.1 and Assumption 4.2 (i) imply the existence of $f_{1}\in R$ such that $f_{i}(x_{k})\geq f_{1}$ for all $k\geq 0$ and $i\in[m]$. It follows by (4.1) and (4.17) that
$\|\gamma_{k}\|\leq(2\max_{i\in[m]}L_{i}+\max_{i\in[m]}(f_{i}(x_{0})-f_{1}))\|s_{k}\|,$
hence,
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq\frac{c_{2}}{\|d_{k}\|_{B_{k}}^{2}},$
where $c_{2}=\frac{(2\max_{i\in[m]}L_{i}+\max_{i\in[m]}(f_{i}(x_{0})-f_{1}))^{2}}{c_{1}}.$
Remark 4.2.
$\beta_{k}$ is the angle between $s_{k}$ and $B_{k}s_{k}$. By Lemma 4.2 and Proposition 4.2, we can obtain that the angles $\beta_{k}$ remain away from 0, for an arbitrary fraction $p$ of the iterates.
Lemma 4.3.
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Then,
$\mathcal{D}(x_{k},d_{k})\leq-\cos\beta_{k}\|d_{k}\|\|d_{SD}^{k}\|.$ (4.18)
Proof.
Since $s_{k}=x_{k+1}-x_{k}=\alpha_{k}d_{k}$, then
$\cos\beta_{k}=\frac{s_{k}^{\top}B_{k}s_{k}}{\|s_{k}\|\|B_{k}s_{k}\|}=\frac{d_{k}^{\top}B_{k}d_{k}}{\|d_{k}\|\|B_{k}d_{k}\|}.$
It follows that
$\mathcal{D}(x_{k},d_{k})$ $\leq-\|d_{k}\|_{B_{k}}^{2}$
$=-\cos\beta_{k}\|d_{k}\|\|B_{k}d_{k}\|$
$=-\cos\beta_{k}\|d_{k}\|\left\|\sum_{i\in[m]}\lambda_{i}^{k}\nabla f_{i}(x_{k})\right\|$
$\leq-\cos\beta_{k}\|d_{k}\|\|d_{SD}^{k}\|.$
This completes the proof.
The following result establishes that the sequence $\{x_{k},d_{k}\}$ satisfies the Zoutendijk condition, which is the basis for establishing the global convergence of Algorithm 1.
Lemma 4.4.
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Suppose that Assumption 4.1 and Assumption 4.2 hold. Then
$\sum_{k\geq 0}\frac{\mathcal{D}(x_{k},d_{k})}{\|d_{k}\|^{2}}<\infty.$ (4.19)
Proof.
The proof is similar to that in [14, Proposition 3.3], we omit it here. ∎
The following theorem establishes the global convergence of Algorithm 1.
Theorem 4.2.
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Suppose that Assumption 4.1 and Assumption 4.2 hold. Then, the following statements hold.
(i) $\liminf\limits_{k\to\infty}\|d_{SD}^{k}\|=0$,
(ii) Every accumulation point of the sequence $\left\{{{x}_{k}}\right\}$ is Pareto critical.
Proof.
Assume that there exists a constant $\varepsilon>0$ such that
$\|d_{SD}^{k}\|\geq\varepsilon,\quad\forall k\in N.$ (4.20)
By the definition of $d_{k}$ and Assumption 4.2 (iii), we get
$-\frac{1}{2}\|d_{k}\|_{B_{k}}^{2}$ $=\min\limits_{d\in R^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}d+\frac{1}{2}}\|d\|_{B_{k}}^{2}$
$\leq\min\limits_{d\in R^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}d+\frac{b}{2}}\|d\|^{2}$
$\leq\frac{1}{b}\min\limits_{d\in R^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}(bd)+\frac{1}{2}}\|bd\|^{2}$
$=-\frac{1}{2b}\|d_{SD}^{k}\|^{2}.$
Set $C_{1}=\frac{c_{1}\varepsilon^{2}}{b}$ and $C_{2}=\frac{c_{2}b}{\varepsilon^{2}}$, it follows by Proposition 4.2 that
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq C_{1}>0,$
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq C_{2}.$
Then, by Lemma 4.3 and Lemma 4.2, there exist a constant $\delta>0$ and a infinite subsequence of indices $\mathcal{K}\subset N$ such that
$-\frac{\mathcal{D}(x_{k},d_{k})}{\|d_{k}\|}\geq\delta\|d_{SD}^{k}\|\geq\delta\varepsilon,\quad\forall k\in\mathcal{K},$
it follows that
$\sum\limits_{k\in N}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}\geq\sum\limits_{k\in\mathcal{K}}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}\geq\sum\limits_{k\in\mathcal{K}}\delta^{2}\varepsilon^{2}=\infty,$
this contradicts the result of Lemma 4.4. Then,
$\liminf\limits_{k\to\infty}\|d_{SD}^{k}\|=0.$
Assumption 4.1 implies there exists a infinite subsequence of indices $K_{1}\subset{N}$ such that $x_{k}\stackrel{{\scriptstyle K_{1}}}{{\longrightarrow}}x^{*}$ and $\|d_{SD}^{k}\|\stackrel{{\scriptstyle K_{1}}}{{\longrightarrow}}0.$ By the Lemma 3.1, $d_{SD}(x)$ is continuous, we can get
$d_{SD}(x^{*})=0.$
Furthermore, $x^{*}$ is Pareto critical. ∎
4.2 Superlinear convergence analysis
In this section, we consider the local superlinear convergence of Algorithm 1 under Assumption 4.3.
Assumption 4.3.
(i) $F$ is twice continuously differentiable.
(ii) For $i\in[m]$, $\{\gamma_{k}\}$ satisfies $\sum\limits_{k>0}\gamma_{k}<\infty.$
(iii) There exists constants $U,L>0$ such that $UI_{n}\preceq\nabla^{2}f_{i}(x)\preceq LI_{n}.$
(iv) $\nabla^{2}f_{i}(x)$ is Lipschitz continuous with constant $M$, $i\in[m].$
(v)For any $x\in R^{n}$, there exists constant $a$ and $b$ such that
$aI_{n}\preceq B(x)\preceq bI_{n}.$
The Assumption 4.3 $(iii)$ implies that $f_{i}(x)$ is strongly convex with modulus $U$ and $\nabla^{2}f_{i}(x)$ is Lipschitz continuous with constant $L$, for $i\in[m]$. In the previous section, we obtain the weak convergence of the sequence $\{x^{k}\}$ generated by Algorithm 1. Under strict convexity, the following theorem demonstrates that the sequence $\{x^{k}\}$ generated by Algorithm 1 strongly converges to the Pareto optimal point $x^{*}.$
Theorem 4.3.
Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Suppose that $f(x)$ is strictly convex, Assumption 4.1 and Assumption 4.2 hold. Then, $\{x_{k}\}$ converges to Pareto optimal point $x^{*}$ of (MOP).
Proof.
The proof is similar to that in [20, Theorem 3.6], we omit it here. ∎
Next, we investigate the R-linear convergence of Algorithm 1.
Lemma 4.5.
Suppose that Assumption 4.3 holds and $\{x_{k}\}$ is the sequence generated by Algorithm 1. Then, for any $k\geq 0$,
(i) $\|x_{k}-x^{*}\|\leq\frac{2}{U}\|d_{SD}^{k}\|,$
(ii) $\|s_{k}\|\geq\frac{1-\sigma_{2}}{L}\cos\beta_{k}\|d_{SD}^{k}\|.$
Proof.
The two assertions can be obtained by using the same arguments as in the proof of [19, Lemma 4.4], we omit them here. ∎
Proposition 4.3.
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Suppose that Assumption 4.3 holds. Then, there exists constant $r$ such that
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq U,$ (4.21)
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq\frac{r^{2}}{U}.$ (4.22)
Proof.
By the definition of $\gamma_{k}$, $s_{k}$ and $m_{k}>0$, we have
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}=\frac{(y_{k}+m_{k}s_{k})^{\top}s_{k}}{\|s_{k}\|^{2}}=\frac{y_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}+m_{k}\geq U.$ (4.23)
Since each objective function is U-strongly convex, for $i\in[m]$ and any $x\in{{R}^{n}}$, ${{f}_{i}}(x)$ has a lower bound $f_{i}^{\min}.$
It follows from the proof of Proposition 4.2 that
$\|\gamma_{k}\|\leq r\|s_{k}\|,$
where $r=2\max\limits_{i\in[m]}L_{i}+\max\limits_{i\in[m]}(f_{i}(x_{0})-f_{i}^{min}).$
Theorem 4.4.
Suppose that Assumption 4.3 holds and $\{x_{k}\}$ is the sequence generated by Algorithm 1. Let $x^{*}$ be defined as Theorem 4.3. Then,
(i) $\{x_{k}\}$ converges R-linearly to $x^{*},$
(ii) $\sum\limits_{k\geq 0}\|x_{k}-x^{*}\|<\infty.$
Proof.
The proof of the theorem is similar to the corresponding discussion in [21, Proposition 4.1], so we omit the proof here. ∎
Remark 4.3.
Using the proof of Theorem 4.4, it can directly deduce that a constant $r\in(0,1)$ exists such that
$0<\sum\limits_{i\in[m]}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))<r^{k},\ \forall k\geq 0,$
where $\lambda_{k}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\dots,\lambda_{m}^{k}\right)}^{T}}$is the solution of (3.9). In the strongly convex case, we can observe that $\eta_{k}>0$ and
$\sum\limits_{i\in[m]}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))\to 0,\ k\to\infty,$
i.e.,
$m_{k}$ $=\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))$
$=\sum_{i=1}^{m}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))\to 0,\ k\to\infty.$
in (4.5). In this case, $B_{k+1}s_{k}\approx y_{k}$. Then (4.8) can be seen as an approximation of the secant equation in [15]. Remarkably, for strongly convex functions, both MFONMO and MQNMO in [15] may exhibit comparable convergence rates. This observation is supported by the numerical experiments presented in the following section.
[…]
Denote the average matrix $\bar{G}_{i}^{k}$ as follows
[…]
Then,
$\bar{G}_{i}^{k}s_{k}=\mu_{i}^{k}.$ (4.24)
Next, we define another two auxiliary functions as follows
$f_{\lambda}(x)=\sum\limits_{i\in[m]}\lambda_{i}^{k}f_{i}(x),$
$G_{\lambda}^{k}=\int_{0}^{1}\sum\limits_{i\in[m]}\lambda_{i}^{k}\nabla^{2}f_{i}(x_{k}+\tau s_{k})d\tau.$
where $\lambda_{k}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\dots,\lambda_{m}^{k}\right)}^{T}}$is the solution of (3.9). Let us define
$\bar{s}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}s_{k},\quad\bar{y}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}y_{k},\quad\bar{\gamma}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\gamma_{k},$
$\bar{B}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}B_{k}\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}},$ (4.25)
$\cos\bar{\beta}_{k}=\frac{\bar{s}_{k}^{\top}\bar{B}_{k}\bar{s}_{k}}{\|\bar{s}_{k}\|\|\bar{B}_{k}\bar{s}_{k}\|},\quad\bar{q}_{k}=\frac{\bar{s}_{k}^{\top}\bar{B}_{k}\bar{s}_{k}}{\bar{s}_{k}^{\top}\bar{s}_{k}},\quad\bar{a}_{k}=\frac{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}},\quad\bar{b}_{k}=\frac{\|\bar{\gamma}_{k}\|^{2}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}.$
By the BFGS-tpye matrix updated fomula (4.6), it can be noted that (4.25) leads to the formula
$\bar{B}_{k+1}=\bar{B}_{k}-\dfrac{\bar{B}_{k}\bar{s}_{k}\bar{s}_{k}^{\top}\bar{B}_{k}}{\bar{s}_{k}^{\top}\bar{B}_{k}\bar{s}_{k}}+\dfrac{\bar{\gamma}_{k}\bar{\gamma}_{k}^{\top}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}.$
Another useful tool used in [22], given that the following function
$\psi(B)=\rm trace(B)-\ln(\det(B)),$
$\psi(\bar{B}_{k+1})=\psi(\bar{B}_{k})+(\bar{b}_{k}-\ln(\bar{a}_{k})-1)+\ln(\cos^{2}\bar{\beta}_{k})+\left[1-\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}+\ln(\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}})\right]$ (4.26)
[…]
Theorem 4.5.
Suppose that Assumption 4.3 holds. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then
$\lim\limits_{k\to\infty}\frac{\|(B_{k}-\nabla^{2}f_{\lambda}(x^{*}))s_{k}\|}{\|s_{k}\|}=0.$
Proof.
By (4.24) and $\lambda_{k}\in{\Delta}_{m}$, we can get
$\bar{G}_{\lambda}^{k}s_{k}=y^{k},$
hence,
$(\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}(x^{*}))s_{k}=y^{k}-\nabla^{2}f_{\lambda}(x^{*})s_{k},\quad k\geq 0.$ (4.27)
It follows by Assumption 4.3 (iv), (4.27) and the definition of $\bar{y}_{k}$, $\bar{s}_{k}$ that
$\|\bar{y}_{k}-\bar{s}_{k}\|$ $=\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}(\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}(x^{*}))\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\bar{s}_{k}\|$
$\leq\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|^{2}\|\bar{s}_{k}\|\|\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}(x^{*})\|$
[…]
$\leq\bar{c}\|\bar{s}_{k}\|\varepsilon_{k},$ (4.28)
where $\bar{c}=M\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|^{2}$, $\varepsilon_{k}=\max\{\|x_{k+1}-x^{*}\|,\|x_{k}-x^{*}\|\}.$ Since
$|\|\bar{y}_{k}\|-\|\bar{s}_{k}\||\leq\|\bar{y}_{k}-\bar{s}_{k}\|,$
by (4.28), we get
$(1-\bar{c}\varepsilon_{k})\|\bar{s}_{k}\|\leq\|\bar{y}_{k}\|\leq(1+\bar{c}\varepsilon_{k})\|\bar{s}_{k}\|,$ (4.29)
$(1-\bar{c}\varepsilon_{k})^{2}\|\bar{s}_{k}\|^{2}-2\bar{y}_{k}^{\top}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}\leq\|\bar{y}_{k}\|^{2}-2\bar{y}_{k}^{\top}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}\leq\bar{c}^{2}\|\bar{s}_{k}\|^{2}\varepsilon_{k}^{2}.$
It follows that
$\frac{\bar{y}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}\geq 1-\bar{c}\varepsilon_{k}.$ (4.30)
Using the Assumption 4.3 (iii), and (4.30) we obtain
$\bar{a}_{k}=\frac{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}=\frac{(\bar{y}_{k}+m_{k}\nabla^{2}f_{\lambda}(x^{*})^{-1}\bar{s}_{k})^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}$ $=\frac{\bar{y}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}+m_{k}\frac{\bar{s}_{k}^{\top}\nabla^{2}f_{\lambda}(x^{*})^{-1}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}$
$\geq 1-\bar{c}\varepsilon_{k}+\frac{m_{k}}{L}$
$\geq 1-\bar{c}\varepsilon_{k}.$ (4.31)
By the L-Lipschitz continuity of $f_{i}$, we have
$\frac{\|y_{k}\|}{\|s_{k}\|}\leq L.$ (4.32)
From (4.21), (4.32) and the right side of (4.29),
$\bar{b}_{k}$ $=\frac{\|\bar{\gamma}_{k}\|^{2}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}$
$=\frac{\|\bar{y}_{k}\|^{2}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}+2m_{k}\frac{\bar{y}_{k}^{\top}\nabla^{2}f_{\lambda}(x^{*})^{-1}\bar{s}_{k}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}+(m_{k})^{2}\frac{\bar{s}_{k}^{\top}\nabla^{2}f_{\lambda}(x^{*})^{-2}\bar{s}_{k}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}$
$\leq\frac{(1+\bar{c}\varepsilon_{k})^{2}}{1-\bar{c}\varepsilon_{k}}+2m_{k}\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|\|y_{k}\|\|\nabla^{2}f_{\lambda}(x^{*})^{-1}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|\|s_{k}\|}{U\|s_{k}\|^{2}}+(m_{k})^{2}\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-2}\|\|\bar{s}_{k}\|^{2}}{U\|s_{k}\|^{2}}$
$\leq 1+\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{k}}{1-\bar{c}\varepsilon_{k}}\varepsilon_{k}+\eta_{1}m_{k}+\eta_{2}m_{k}^{2},$ (4.33)
where $\eta_{1}=2\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|\|\nabla^{2}f_{\lambda}(x^{*})^{-1}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|L}{U}$, $\eta_{2}=\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-2}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|^{2}}{U}.$ Using (4.26) and (4.31),
$\psi(\bar{B}_{k+1})\leq\psi(\bar{B}_{k})$ $+\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{k}}{1-\bar{c}\varepsilon_{k}}\varepsilon_{k}+\eta_{1}m_{k}+\eta_{2}m_{k}^{2}-\ln(1-\bar{c}\varepsilon_{k})\right)$
$+\ln(\cos^{2}\bar{\beta}_{k})+\left[1-\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}+\ln\left(\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}\right)\right],$
By adding the above inequality from 0 to k, then we obtain
$0<\psi(\bar{B}_{k+1})\leq\psi(\bar{B}_{1})$ $+\sum_{j=1}^{k}\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{j}}{1-\bar{c}\varepsilon_{j}}\varepsilon_{j}+\eta_{1}m_{j}+\eta_{2}m_{j}^{2}-\ln(1-\bar{c}\varepsilon_{j})\right)$
$+\left[1-\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}+\ln\left(\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}\right)\right]+\ln\left(\cos^{2}\bar{\beta}_{j}\right),$ (4.34)
From Theorem 4.4, there exists $k_{0}\in N$ such that
$\bar{c}\varepsilon_{k}<\frac{1}{2},\ \forall k>k_{0},$
then, $\ln(1-\bar{c}\varepsilon_{k})\geq-2\bar{c}\varepsilon_{k}.$ It follows by Assumption 4.3 (ii) that
$\sum\limits_{j>k_{0}}\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{j}}{1-\bar{c}\varepsilon_{j}}\varepsilon_{j}+\eta_{1}m_{j}+\eta_{2}m_{j}^{2}-\ln(1-\bar{c}\varepsilon_{j})\right)$ $<\sum\limits_{j>k_{0}}9\bar{c}\varepsilon_{j}+\sum\limits_{j>k_{0}}(\eta_{1}m_{j}+\eta_{2}m_{j}^{2})$
$<\sum_{j=1}^{\infty}9\bar{c}\varepsilon_{j}+\sum_{j=1}^{\infty}(\eta_{1}m_{j}+\eta_{2}m_{j}^{2})$
<<$\infty$.< ∞ . (4.35)
By (4.35) and (4.34), we get
$\sum_{j=1}^{\infty}\left(-\left[1-\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}+\ln\left(\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}\right)\right]-\ln\left(\cos^{2}\bar{\beta}_{j}\right)\right)<\infty.$
Since $-\ln(\cos^{2}\bar{\beta}_{j})>0$ ,for all $j>k_{0}$, and $h(t)=1-t+lnt\leq 0$ when $t>0$, hence, we get
$-\left[1-\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}+\ln\left(\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}\right)\right]\geq 0,$
$\lim\limits_{k\to\infty}\ln\left(\frac{1}{\cos^{2}\bar{\beta}_{k}}\right)=0,\quad\lim\limits_{k\to\infty}\left[1-\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}+\ln\left(\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}\right)\right]=0.$
The above relations imply that
$\lim\limits_{k\to\infty}\ln\left(\cos^{2}\bar{\beta}_{j}\right)=1,\quad\lim\limits_{k\to\infty}\bar{q}_{k}=1.$
Now,
$\lim\limits_{k\to\infty}\frac{\|(B_{k}-\nabla^{2}f_{\lambda}(x^{*}))s_{k}\|^{2}}{\|s_{k}\|^{2}}$ $=\lim\limits_{k\to\infty}\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}(B_{k}-\nabla^{2}f_{\lambda}(x^{*}))\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\bar{s}_{k}\|^{2}}{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}s_{k}\|^{2}}$
$=\lim\limits_{k\to\infty}\frac{\|(\bar{B}_{k}-I_{n})\bar{s}_{k}\|^{2}}{\|\bar{s}_{k}\|^{2}}$
$=\lim\limits_{k\to\infty}\frac{\|\bar{B}_{k}\bar{s}_{k}\|^{2}-2\bar{s}_{k}^{\top}\bar{B}_{k}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}}{\|\bar{s}_{k}\|^{2}}$
$=\lim\limits_{k\to\infty}\left[\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}-2\bar{q}_{k}+1\right]$
$=0$
Theorem 4.6.
Suppose that Assumptions 4.3 hold. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then,
(i) (4.1) and (4.2) hold with $\alpha_{k}=1$ for sufficiently large k.
(ii) $\{x_{k}\}$ converge to $x^{*}$ at a superlinear rate.
Proof.
The proof is similar to that in [20, Theorem4.3], we omit it here. ∎
5 Numerical experiments
In this section, we present some numerical experiments to verify the efficiency of MFQNMO. We compare MFQNMO with QNMO [9] with Wolfe line search and MQNMO [15] with Wolfe line search. The selected test problems include convex and nonconvex problems. Our evaluation criteria include the method’s effectiveness in finding optimal solutions and its ability to approximate real Pareto surfaces accurately. These experiments were implemented using Python 3.9.
For MQNMO and MFQNMO methods, we used the Frank-Wolf gradient method to solve the dual problem (3.9). For QNMO, (3.4) is solved using scipy. optimize, which is an optimization Solver. All test problems and main characteristics are listed in Table 1. Columns ‘n’ and ‘m’ indicate the number of variables and objectives, respectively. The ‘Convex’ column indicates whether the problem is convex or not. ’Y’ stands for Yes. ’N’ stands for NO. Problems are solved under box constraints ${{x}_{L}}\leq x\leq{{x}_{U}}.$ The initial points are randomly selected within these box constraints ${{x}_{L}}\leq x\leq{{x}_{U}}.$ We used the termination criterion $\left|\theta\left(x\right)\right|\leq{{10}^{-8}}$ for all test problems. The maximum number of iterations is set to 500. To obtain the stepsize that satisfies the Wolfe conditions (4.1) and (4.2), we used the Wolf line search algorithm proposed in [21]. We used $\sigma_{1}={10}^{-4}$, $\sigma_{2}=0.1$, and set $B_{0}=I_{n}$, for all $i\in[m]$. We perform 200 computations using the same initial points for all methods.
Table 1: The description of all problems used in numerical experiments
problem $n$ $m$ ${{x}_{L}}$ ${{x}_{U}}$ convex Reference
SD 4 2 $(1,-\sqrt{2},-\sqrt{2},1)$ $(3,3,3,3)$ Y [23]
PNR 2 2 $(-2,-2)$ $(2,2)$ Y [24]
JOS1a 50 2 $(-2,-2)$ $(2,2)$ Y [25]
JOS1b 100 2 $(-2,-2)$ $(2,2)$ Y [25]
DGO1 1 2 $-10$ $13$ N [26]
DGO2 1 2 $-9$ $9$ Y [26]
Lov1 2 2 $(-10,-10)$ $(10,10)$ Y [27]
Lov2 2 2 $(-0.75,-0.75)$ $(0.75,0.75)$ N [27]
Lov3 2 2 $(-20,-20)$ $(20,20)$ N [27]
Lov4 2 2 $(-20,-20)$ $(20,20)$ N [27]
SK1 1 2 $-100$ $100$ N [26]
BK1 2 2 $(-5,-5)$ $(10,10)$ Y [26]
SLCDT1 2 2 $(-1.5,-1.5)$ $(1.5,1.5)$ N [28]
MOP1 1 2 $-{10}^{5}$ ${10}^{5}$ Y [26]
MOP2 2 2 $(-4,-4)$ $(4,4)$ N [26]
LDTZ 3 3 $(0,0,0)$ $(1,1,1)$ N [29]
Hil1 2 2 $(0,0)$ $(1,1)$ N [30]
AP2 1 2 $-100$ $100$ Y [15]
AP3 2 2 (100,100))(-100,-100))( - 100 , - 100 ) ) $(100,100)$ N [15]
FF1 2 2 $(-1,-1)$ $(1,1)$ N [26]
KW2 2 2 $(-3,-3)$ $(3,3)$ N [31]
MHHM1 1 3 00 $1$ Y [26]
MHHM2 2 3 $(0,0)$ $(1,1)$ Y [26]
Table 2: Number of average iterations (iter), average CPU time (time(ms)), number of average function evaluations (feval), and number of failure points (NF) of MFQNMO, QNMO, and MQNMO.
Problem QNMO MFQNMO MQNMO
iter time feval NF iter time feval NF iter time feval NF
SD 13.49 486.47 15.54 1 7.35 2.55 18.51 1 11.17 5.53 43.54 1
PNR 2.56 118.66 5.48 0 7.87 2.81 28.31 0 8.12 3.84 54.67 0
JOS1a 3.01 270.14 7.39 0 5.96 3.84 32.89 0 3.05 5.38 24.51 0
JOS1b 2.89 12725.46 7.39 0 7.65 19.38 10.05 0 11.23 43.337 103.01 0
DGO1 175.51 1906.25 5058.35 70 1.39 0.64 4.14 0 1.44 0.72 4.37 0
DGO2 3.59 52.34 9.56 0 5.40 1.64 17.41 0 70.35 31.75 409.96 0
Lov1 268.91 3969.94 9213.41 107 4.27 1.45 12.72 0 4.30 1.72 15.04 0
Lov2 30.43 448.92 77.89 146 12.36 8.41 31.49 124 29.70 19.44 107.27 124
Lov3 78.77 1462.87 3400.81 30 11.86 6.27 29.42 3 24.13 9.81 71.69 3
Lov4 248.53 4516.95 12193.75 101 1.30 0.57 4.04 0 1.52 0.71 4.85 0
SK1 119.17 488.25 2747.57 69 2.36 1.65 26.54 0 2.66 1.93 46 0
BK1 98.41 1085.38 100.41 39 1 0.13 3 0 1 0.16 3 0
SLCDT1 F F F 200 3.35 0.64 8.51 0 4.18 1.38 11.28 0
MOP1 4.66 237.83 146.13 47 1.09 0.21 4.89 18 1 0.15 3 0
MOP2 89.24 1489.39 34344.11 35 3.67 1.53 11.41 0 7.18 2.28 121.99 1
LDTZ F F F 200 18.02 87.05 119.06 18 F F F 200
Hil1 258.88 7409.41 6274.12 110 11.71 6.41 62.52 16 98.39 186.89 529.98 53
AP2 1.04 35.95 3.17 1 0.985 0.41 2.95 0 0.985 0.46 2.95 0
AP3 147.91 3363.81 5353.84 115 22.36 10.41 109.45 0 248.21 204.68 3858.64 77
FF1 245.17 3394.94 10089.34 98 32.05 15.97 190.12 0 16.12 8.94 81.77 0
KW2 92.97 1697.45 298.53 103 13.81 7.98 51.01 24 99.81 69.37 314.52 32
MHHM1 0.88 26.87 2.82 0 0.88 0.40 2.65 0 0.88 0.42 2.65 0
MHHM2 245.54 4752.67 9049.92 99 6.21 43.96 206.47 31 78.34 1356.45 3508.36 31
In Table 2, the average values obtained from these runs are recorded to determine the number of iterations (iter), function evaluations (feval), and CPU time (time(ms)). The number of points that did not converge in 500 iterations among the 200 initial points is recorded in column ‘NF’ for MFQNMO, QNMO, and MQNMO. We denote ’Failed’ as ’F,’ which indicates that the corresponding method fails to solve the test problem under the stopping criterion. For large-scale problems, such as JOS, the QNMO method’s CPU time consumption is substantial although it requires the fewest iterations. For all nonconvex test problems in Table 1, the QNMO method exhibited poor performance and the MQNMO method surpasses the MFQNMO method only for the FF1 test problem. For certain problems in Table 1, the experimental efficiency of the MFQNMO method and the MQNMO method are equal, such as BK1, MOP1, MHHM1, and DGO1. We mentioned the possible reasons for this phenomenon in Remark 4.3. As can be seen from Table 2, the MFQNMO method is more efficient and competitive.
Figure 1 shows the Pareto frontier obtained using the MFQNMO method for the convex problem listed in Table 1. Figure 2 shows the Pareto frontier obtained using the MFQNMO method for the nonconvex problem listed in Table 1. For the chosen test problems, it can be observed that the MFQNMO method was able to satisfactorily estimate the Pareto frontiers when considering evenly distributed starting points.
(a) Lov1
(b) PNR
(c) DGO2
(d) BK1
Figure 1: The approximated Pareto frontiers generated by MFQNMO for the convex problems Lov1, PNR, DGO2, and BK1.
(a) Lov3
(b) Lov4
(c) DGO1
(d) FF1
(e) Hil
(f) KW2
(g) LDTZ
(h) MOP2
(i) SD
(j) SK1
(k) SLCDT1
Figure 2: The approximated Pareto frontiers generated by MFQNMO for the nonconvex problems Lov3, Lov4, DGO1, FF1, Hil, KW2, LTDZ, MOP2, SD, SK1, and SLDTZ1.
6 Conclusion
We propose a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems (MFQNMO), in which the BFGS-type formula is updated by using the function information. To obtain the global convergence property, we only need the smoothness of the objective function. Compared with [20], our work has the following advantages.
(1) Instead of approximating the Hessian matrix of all objective functions, we use a common matrix to capture the curvature information of the objective function, which balances the effectiveness of the algorithm and the computational cost.
(2) For nonconvex functions, most versions of quasi-Newton methods may not converge, so the MFQNMO method proposed in this paper for nonconvex unconstrained problems is meaningful.
(3) Compared with [20], the update of the BFGS-type formula does not depend on the selection of parameters, but only needs to use the information of past iterations.
In our future work, we will consider applying the method to nonconvex constrained optimization problems.